OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (nikato):

So let x= 1st integer Let x+1=2nd integer And x+2=3rd integer And since u know the sum of these consecutive integer is 267 You can write this equation x+x+1+x+2=267 So now can you solve for x from here?

OpenStudy (anonymous):

X = 90?

OpenStudy (anonymous):

88

OpenStudy (nikato):

Combine like terms 3x+3=267 -3. -3 -------------- 3x=264 --- ---- 3. 3 x=88

OpenStudy (nikato):

But since your question ask for the largest integer. It would be 88+2=90 Because our largest integer was x+2 So 90 is right

OpenStudy (anonymous):

Wow okay cool thank you both so much :)

OpenStudy (nikato):

Your welcome!